Х dx 4 1 +X

Integrate using partial fractions.
Answer #1

we are given

\int \frac{x}{1+x^4}dx

we can also write as

\int \frac{x}{1+(x^2)^2}dx

we can use u-subs

LE = n

du=2xdx

we get

=\int \frac{x}{1+u^2}\frac{du}{2x}

=\frac{1}{2}\int \frac{1}{1+u^2}du

=\frac{1}{2}tan^{-1}(u)+C

now, we can plug back u=x^2

=\frac{1}{2}tan^{-1}(x^2)+C............Answer